# Improving GFlowNets for Text-to-Image Diffusion Alignment

### v1

@@ B.1 Experimental details
Regarding training hyperparameters, we follow the DDPO github repository implementation and describe them below for completeness. We use classifier-free guidance (Ho & Salimans, 2022, CFG) with guidance weight being $5$. We use a $50$-step DDIM schedule. We use NVIDIA 8×8$\times$8 ×A100 80GB GPUs for each task, and use a batch size of $8$ per single GPU. We do $4$ step gradient accumulation, which makes the essential batch size to be $256$. For each “epoch”, we sample $512$ trajectories during the rollout phase and perform $8$ optimization steps during the training phase. We train for $100$ epochs. We use a $3\times 10^{-4}$ learning rate for both the diffusion model and the flow function model without further tuning. We use the AdamW optimizer and gradient clip with the norm being $1$. We set $\epsilon=1\times 10^{-4}$ in Equation 14. We use bfloat16161616 precision.
@@
The GFlowNet framework requires the reward function to be always non-negative, so we just take the exponential of the reward to be used as the GFlowNet reward. We also set the reward exponential to $\beta=100$ (i.e., setting the distribution temperature to be $1/100$). Therefore, $\log R(\cdot)=\beta R_{\text{original}}(\cdot)$. Note that in GFlowNet training practice, we only need to use the logarithm of the reward rather than the original reward value. We linearly anneal $\beta$ from 00 to its maximal value in the first half of the training. We found that this almost does not change the final result but is helpful for training stability.For DAG-KL, we put the final $\beta$ coefficient on the KL gradient term. We also find using a KL regularization $\mathcal{D}_{\mathrm{KL}}\left(p_{\boldsymbol{\theta}}(\mathbf{x}_{t-1}|\mathbf{x}_{t})\|p_{{\boldsymbol{\theta}}_{\text{old}}}(\mathbf{x}_{t-1}|\mathbf{x}_{t})\right)$ to be helpful for stability (this is also mentioned in  Fan et al. (2023)). In practice, it is essentially adding a $\ell_{2}$ regularization term on the output of the U-Net after CFG between the current model and previous rollout model. We simply use a coefficient $1$ on this term without further tuning.
@@
We use Stable Diffusion v1.5 as base model and use LoRA for post-training following Black et al. (2023). For the architecture of the state flow function, we take a similar structure to the downsample part of the U-Net. The implementation is based on the hugging face diffusers package. We use $3$ “CrossAttnDownBlock2D” blocks and $1$ “DownBlock2D” and do downsampling on all of them. We set the layers per block to be $1$, and set their block out channels to be $64,128,256,256$. We use a final average pooling layer with kernel and stride size $4$ to output a scalar given inputs including latent image, time step, and prompt embedding. We do not report diversity metric as in previous GFlowNet literature, as the average pairwise Euclidean distance in high dimensional space ($64\times 64\times 4>10,000$ dim.) is not a meaningful metric.
B.2 CIFAR-10 toy example

### v2

@@ D.2 Experimental details
Regarding training hyperparameters, we follow the DDPO github repository implementation and describe them below for completeness. We use classifier-free guidance (ho2022classifierfree, CFG) with guidance weight being $5$. We use a $50$-step DDIM schedule. We use NVIDIA 8×8$\times$8 ×A100 80GB GPUs for each task, and use a batch size of $8$ per single GPU. We do $4$ step gradient accumulation, which makes the essential batch size to be $256$. For each “epoch”, we sample $512$ trajectories during the rollout phase and perform $8$ optimization steps during the training phase. We train for $100$ epochs. We use a $3\times 10^{-4}$ learning rate for both the diffusion model and the flow function model without further tuning. We use the AdamW optimizer and gradient clip with the norm being $1$. We set $\epsilon=1\times 10^{-4}$ in Equation 11. We use bfloat16161616 precision.
@@
The GFlowNet framework requires the reward function to be always non-negative, so we just take the exponential of the reward to be used as the GFlowNet reward. We also set the reward exponential to $\beta=100$ (i.e., setting the distribution temperature to be $1/100$). Therefore, $\log R(\cdot)=\beta R_{\text{original}}(\cdot)$. Note that in GFlowNet training practice, we only need to use the logarithm of the reward rather than the original reward value. We linearly anneal $\beta$ from 00 to its maximal value in the first half of the training. We found that this almost does not change the final result but is helpful for training stability.For DAG-KL, we put the final $\beta$ coefficient on the KL gradient term. We also find using a KL regularization $\mathcal{D}_{\mathrm{KL}}\left(p_{\boldsymbol{\theta}}(\mathbf{x}_{t-1}|\mathbf{x}_{t})\|p_{{\boldsymbol{\theta}}_{\text{old}}}(\mathbf{x}_{t-1}|\mathbf{x}_{t})\right)$ to be helpful for stability (this is also mentioned in  Fan2023DPOKRL). In practice, it is essentially adding a $\ell_{2}$ regularization term on the output of the U-Net after CFG between the current model and previous rollout model. We simply use a coefficient $1$ on this term without further tuning.
@@
We use Stable Diffusion v1.5 as base model and use LoRA for post-training following Black2023TrainingDM. For the architecture of the state flow function, we take a similar structure to the downsample part of the U-Net. The implementation is based on the hugging face diffusers package. We use $3$ “CrossAttnDownBlock2D” blocks and $1$ “DownBlock2D” and do downsampling on all of them. We set the layers per block to be $1$, and set their block out channels to be $64,128,256,256$. We use a final average pooling layer with kernel and stride size $4$ to output a scalar given inputs including latent image, time step, and prompt embedding. We do not report diversity metric as in previous GFlowNet literature, as the average pairwise Euclidean distance in high dimensional space ($64\times 64\times 4>10,000$ dim.) is not a meaningful metric.
D.3 CIFAR-10 toy example